Two scanners are used to scan a large document. "Scanner A" can scan the whole document in 3 hours; "Scanner B" is slower and can scan the same document in 5 hours. Since BOTH scanners are being used, how long will it take for the entire document to be scanned?
Answer by TimothyLamb(4379)

scan rate of both scanners together:
r = a + b
where:
a = scan rate of fast scanner = 1/3 = 0.3333 documents/hr
b = scan rate of slow scanner = 1/5 = 0.2 documents/hr
---
r = a + b
r = 1/3 + 1/5
r = 5/15 + 3/15
r = 8/15
r = 0.5333 documents/hr
---
scan speed of both scanners together:
s = 1/r
s = 1/0.5333
s = 1.875 hr/document
s = 1.875 hr/document * 60 min/hr = 112.5 min
---
Answer:
both scanners together can scan the document in 1 hour, 52 minutes, and 30 seconds
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php